# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 9.4.3 The adjoint SU(2) gauge Higgs model with four flavors
Let us now focus on the model with $N_{c}=2$ and $N_{f}=4$, which is expected to describe some aspects of the optimal doping criticality in cuprate high-$T_{c}$ superconductors [96]. Since $N_{f}>N_{c}^{2}-1$, there are various Higgs phases characterized by different global and gauge symmetry-breaking patterns. For finite $\beta$ the two Higgs phases are divided by a transition line that ends at $v=0$, $\beta=\infty$, and which is expected to be generically of first order, as it is the boundary of two different ordered phases.
Figure 23: Sketches of the phase diagram for $N_{c}=2$, $N_{f}=4$ [98], at fixed $\gamma=0$ (left) and at fixed $v=24$ (right). Left panel: thick lines denote first-order transitions, while the thin line corresponds to continuous transitions; the shaded point is a first-order meeting point; the filled black point that separates first-order from continuous transitions occurs at $v=v^{*}$ with $126<v^{*}<$ 12. Right panel: the ${\mathbb{Z}}_{2}$ gauge transition line starts at $\kappa_{c}\approx 0.761$, $\beta=\infty$. The multicritical point, where the three transition lines meet, satisfies $21<\kappa<$ 2. In each Higgs phase we also report the gauge-symmetry-breaking pattern associated with the transition that divides the Higgs phase from the disordered phase.
@@
For larger values of $\gamma$ it is possible to have topological transitions in the positive-$v$ region. Indeed, one may have two different positive-$v$ Higgs phases, which differ for the behavior of the topological modes associated with the gauge-group center ${\mathbb{Z}}_{2}$ [96, 259]. Some numerical evidence of the existence of such topological transitions in the Higgs phase has been reported in Ref. [98], looking at the scaling behavior of the energy cumulants, as described in A.3.4.
@@
To summarize, the phase diagram shown in the right panel of Fig. 23 appears to be characterized by three phases, a small-$\beta$ disordered phase, and two large-$\beta$ Higgs phases, which are distinguished by the behavior of gauge-group center modes. These phases are separated by three transition lines: (i) a disordered-Higgs transition line for small $\kappa$, which appears to be continuous; (ii) a disordered-Higgs transition line for large $\kappa$, which is of first order; (iii) a continuous ${\mathbb{Z}}_{2}$ gauge (Ising) transition line, which separates the two low-temperature Higgs phases.
@@
As a final remark, we note that the identification of continuous transitions associated with the stable charged FP of the RG flow of the adjoint NAHFT, see Sec.8.3, remains an open issue. We believe that the phase diagram and critical behaviors of SU($N_{c}$) gauge LNAH models with adjoint scalar fields call for further numerical investigations, to achieve a satisfactory characterization of their main features.
@@ 9.5 Lattice SO($N_{c}$) gauge models with multiflavor scalar fields
One may again consider the unit-length limit of the scalar field, formally obtained by taking the limit $r,u\to\infty$ keeping $r/u=-1$ fixed, leading to the constraint ${\rm Tr}\,\varphi_{\bm{x}}^{t}\varphi_{\bm{x}}=1$. In the one-flavor case, $N_{f}=1$, the phase diagram of the lattice SO($N_{c}$) gauge model (114) is expected to show only one phase. This can be easily verified for $\gamma=0$, where the model becomes trivial in the unitary gauge, and cannot have any phase transition. For $N_{c}=2$ the gauge group is equivalent to the Abelian group U(1), and the model is equivalent to LAHMs with compact gauge variables discussed in Sec. 6. New critical behaviors are expected only for the non-Abelian cases with $N_{c}\geq 3$. The natural order parameter for the breaking of the SO($N_{f}$) global symmetry group is the gauge-invariant real traceless and symmetric bilinear operator
@@
The field-theoretical $\varepsilon$-expansion analysis based on the NAHFT (94) predicts the existence of charged universality classes for a sufficiently large number $N_{f}$ of components, controlled by the stable charged FP appearing for positive values of $v$. No stable FPs emerge for $v<0$, at least close to four dimensions. However, the statistical system may still develop other critical behaviors described by appropriate LGW field theories, in which the gauge correlations are not critical, such as those expected in the strong-coupling ($\gamma\to 0$) gauge limit.
An analysis of the phase diagram of the model with $v=0$ was reported in Ref. [376]. For $N_{f}\geq 2$ the phase diagram is characterized by two phases: a low-temperature phase in which the order parameter $R_{\bm{x}}^{fg}$ defined in Eq. (116) condenses, and a high-temperature disordered phase. The two phases are separated by a transition line, where the SO$(N_{f})$ symmetry is broken. The gauge parameter $\gamma$, corresponding to the inverse gauge coupling, is not expected to play any particular role, at least for sufficiently small values. This scenario has been checked numerically for several values of $\gamma$. Along the disorder-order transition line only the correlations of the gauge-invariant operator $R_{\bm{x}}^{fg}$ are critical, while gauge modes are not critical. This is consistent with the predictions of the continuum gauge model (94): since stable FPs exist only for $N_{f}\gg 1$, one expects only uncharged transitions for small $N_{f}$. Since the gauge modes do not show critical behaviors, one may predict the critical behavior by using an effective O($N_{f}$)-invariant LGW theory based on a gauge-invariant order parameter associated with the rank-two symmetric real traceless tensor $R_{\bm{x}}^{ab}$. This analysis is reported in Sec. 3.2.2. It predicts first-order transitions for $N_{f}\geq 3$ and continuous transitions for $N_{f}=2$, which belong to the XY universality class for any $N_{c}\geq 3$.
@@
Note that charged continuous transitions are possible for large values of $N_{f}$ according to the RG flow of the SO($N_{c}$) gauge NAHFT, see Sec. 8.2, because of the presence of a stable large-$N_{f}$ charged FP [273, 376, 155]. Like SU($N_{c}$) gauge LNAH models with scalar fields in the fundamental representation, continuous charged transitions are expected to appear for positive values of $v$, sufficiently large values of $\gamma$, and a sufficiently large number $N_{f}$ of components.
10 Two-dimensional lattice gauge models with multiflavor scalar matter
@@
In the previous sections we have mostly addressed 3D Abelian and non-Abelian Higgs models, to investigate the role of the gauge and global symmetries in determining their main thermodynamic properties, such as the Higgs phases, the nature of the phase transitions, and the continuum limits associated with their critical behaviors. Analogous issues can be addressed in lower-dimensional models.
@@
In two dimensions the phase diagram of lattice models with short-range interactions (such as the nearest-neighbor models considered in the previous sections) and continuous global symmetries are generally characterized by a single disordered phase. Indeed, according to the Mermin-Wagner theorem [410, 411], they cannot have magnetized phases characterized by the condensation of an order parameter, and therefore they do not undergo phase transitions driven by the spontaneous breaking of the global symmetry.
@@
However, in the zero-temperature limit they can develop asymptotic critical behavior, whose universal features are determined by the symmetries of the system and are generally associated with 2D nonlinear $\sigma$ models defined on symmetric spaces, see, e.g., Refs. [412, 2, 95]. The O($N$) $\sigma$ model with $N\geq 3$ and the CPN-1 model with $N\geq 2$ are paradigmatic models that show this type of behavior. Indeed, the lattice O($N$) model behaves as the nonlinear $\sigma$ model defined on the O($N$)/O($N-1$) symmetric space, while the CPN-1 lattice model is related to the nonlinear $\sigma$ model defined on the U($N$)/[U(1)$\times$×U($N-1$)] symmetric space, see, e.g., Ref. [2]. In this type of models, the correlation functions in the thermodynamic limit are characterized by a length scale $\xi$ that diverges as $T^{p}e^{c/T}$ for $T\to 0$, where $p$ and $c$ are universal coefficients that only depend on the universality class and that can be determined in perturbation theory (the coefficient $c$ also depends on the normalization of $T$, which can be uniquely fixed perturbatively), see A.5 for details.
@@
In this section we discuss how the interplay of global and gauge symmetries determine the effective low-energy field theory realized in the zero-temperature continuum limit (critical behavior) of 2D lattice gauge models using the standard Wilson formulation of gauge variables [100, 3].
@@ 10.1 General conjecture for the zero-temperature continuum limit
In the case of models characterized by both global and gauge symmetries, the asymptotic zero-temperature critical behavior, realizing the continuum limit, is expected to arise from the interplay between the two different symmetries. For the purpose of understanding which features are relevant and which continuum limits are effectively realized, several 2D lattice Abelian and non-Abelian gauge models have been investigated [416, 417, 221, 418, 409, 419, 420]. The results of these analyses have led to the following general conjecture:
@@
2D field-theoretical models defined on symmetric space have been largely investigated, see, e.g., Ref. [2], and several high-order perturbative computations, in particular of their $\beta$ functions, have been reported in the literature [421, 422, 423, 424, 425, 426, 280, 412, 273, 427, 428, 429, 430, 431, 432, 433, 434, 435, 436, 2, 95]. The O($N$) $\sigma$ and the CPN-1 models are the simplest examples. Other examples of lattice gauge models with critical behavior corresponding to a symmetric-space field theory will be reported in the following sections.
10.2 Continuum limits of multiflavor lattice Abelian and non-Abelian gauge theories
We now summarize the main results obtained for several 2D multiflavor lattice Abelian and non-Abelian gauge models [416, 417, 418, 409, 419, 420]. They support the above conjecture on the nature of their continuum zero-temperature limit. The analysis of the zero-temperature behavior requires the determination of the relevant low-energy degrees of freedom, which correspond to the minimum configurations of the Hamiltonian, and of the residual global symmetry, which essentially determines the appropriate corresponding field-theoretical $\sigma$ model. These analyses are then supplemented by MC nonperturbative studies based on FSS analyses of the numerical data.
10.2.1 Two-dimensional multiflavor Abelian U(1) gauge models
@@
For some particular values of $\gamma$ model (117) becomes equivalent to simpler well-known models. In the limit $\gamma\to\infty$, the gauge link variables are all equal to one modulo gauge transformations, so model (117) becomes equivalent to the standard 2D O($2N$)-symmetric vector model, whose properties are reviewed in A.5. For $\gamma=0$ the AH model (117) is a particular lattice formulation of the CPN-1 model [2, 204, 205, 206, 431, 432]4949492D CPN-1 models have been much studied in the literature, both analytically and numerically, because they provide a theoretical laboratory to understand some of the mechanisms of quantum field theories of fundamental interactions. In particular, they share some notable features with QCD, the theory of the hadronic strong interactions, such as the asymptotic freedom and the so-called $\theta$ dependence related to topology, see, e.g., Refs. [425, 437, 426, 430, 438, 439]. For both $\gamma=0$ and $\gamma=\infty$, the lattice model shows a universal critical behavior for $\beta\to\infty$. In this limit, the correlation length $\xi$ increases exponentially, $\xi\sim\beta^{-p}e^{c\beta}$. Of course, the critical behaviors for $\gamma=0$ and $\gamma\to\infty$ differ, belonging to the universality class of the 2D CPN-1 $\sigma$ model for $\gamma=0$ and to that of the 2D O($2N$) $\sigma$ model for $\gamma\to\infty$.
As in three-dimensional case, see Sec. 6, the zero-temperature critical behavior can be investigated by analyzing the correlations of the gauge-invariant bilinear $Q_{\bm{x}}^{ab}=\bar{z}_{\bm{x}}^{a}z_{\bm{x}}^{b}-\delta^{ab}/N$. The FSS analyses of the MC results reported in Ref. [416] show that the asymptotic behavior is independent of $\gamma$, and corresponds to that found for $\gamma=0$. Therefore, the zero-temperature critical behavior of the 2D lattice U(1) model with an $N$-component complex scalar field belongs to the universality class of the 2D CPN-1 model, thus corresponding to the symmetric space U($N$)/[U(1)$\times$×U($N-1$)]. The independence of the critical behavior on $\gamma$ is related to the subleading behavior of the gauge modes.
10.2.2 Two-dimensional multiflavor SU($N_{c}$) gauge models
We now consider 2D non-Abelian SU($N_{c}$) models defined on square lattices, with multiflavor matter fields in the fundamental and adjoint representations, analogous to those defined in Secs. 9.1.1 and 9.4.1 in the unit-length limit, respectively.
@@ 10.2.2 Two-dimensional multiflavor SU($N_{c}$) gauge models
Figure 24: (Adapted from Ref. [417]) We show numerical results for the 2D LNAH models with $N_{f}=2$, $N_{c}=3$ (left) and $N_{c}=4$ (right), and $\gamma=0$. We plot the Binder parameter $U_{P}$ associated with the operator $P_{\bm{x}}^{fg}$, defined in Eq. (105), versus $\xi_{P}/L$, where $\xi_{P}$ is the second-moment correlation length associated with the correlation function $G_{P}$, defined as in Eq. 145. In both panels, with increasing $L$, data approach the solid line representing the universal scaling curve of the 2D CP1 model (see Ref. [416] for a parameterization), associated with the symmetric space that has the same global symmetry, thus supporting the conjecture reported in Sec. 10.1. As discussed in Sec. A.3.3), this comparison does not involve any nonuniversal normalization.
10.2.3 Two-dimensional multiflavor SO($N_{c}$) gauge models

### v3

@@ 9.4.3 The adjoint SU(2) gauge Higgs model with four flavors
Let us now focus on the model with $N_{c}=2$ and $N_{f}=4$, which is expected to describe some aspects of the optimal doping criticality in cuprate high-$T_{c}$ superconductors [97]. Since $N_{f}>N_{c}^{2}-1$, there are several Higgs phases characterized by different global and gauge symmetry-breaking patterns. For finite $\beta$ the two Higgs phases are divided by a transition line that ends at $v=0$, $\beta=\infty$, which is expected to be generically of first order, as it is the boundary of two different ordered phases.
@@
Figure 23: Sketches of the phase diagram of 3D SU(2) gauge LNAH models with four scalar fields in the adjoint representation [99], at fixed $\gamma=0$ (left) and at fixed $v=24$ (right). Thick lines denote first-order transitions, while thin lines correspond to continuous transitions. Left panel: the shaded point is a first-order intersection point; the filled black point that separates first-order from continuous transitions occurs at $v=v^{*}$ with $126<v^{*}<$ 12. Right panel: the ${\mathbb{Z}}_{2}$-gauge transition line (dashed line) starts at $\kappa_{c}\approx 0.761$, $\beta=\infty$ and ends at a multicritical point (shaded point) at $\kappa=\kappa_{mc}$ with $21<\kappa_{mc}<$ 2. In each Higgs phase we also report the gauge-symmetry-breaking pattern associated with the transition that divides the Higgs phase from the disordered phase.
@@
For larger values of $\gamma$ it is possible to have topological transitions in the positive-$v$ region. Indeed, one may have two different positive-$v$ Higgs phases, which differ for the behavior of the topological modes associated with the gauge-group center ${\mathbb{Z}}_{2}$ [97, 289]. Numerical evidence of the existence of such topological transitions in the Higgs phase has been reported in Ref. [99], looking at the scaling behavior of the energy cumulants, as described in A.3.2.
@@
To summarize, the phase diagram shown in the right panel of Fig. 23 appears to be characterized by three phases, a small-$\beta$ disordered phase, and two large-$\beta$ Higgs phases, which are distinguished by the behavior of gauge-group center modes. These phases are separated by three transition lines: (i) a disordered-Higgs transition line for small $\kappa$, which appears to be continuous; (ii) a disordered-Higgs transition line for large $\kappa$, which is of first order; (iii) a continuous ${\mathbb{Z}}_{2}$ gauge transition line, which separates the two low-temperature Higgs phases.
@@
As a final remark, we note that the identification of continuous transitions associated with the stable charged FP of the RG flow of the adjoint NAHFT, see Sec.8.3, remains an open issue. We believe that further numerical investigations of the phase diagram and of the critical behaviors of SU($N_{c}$) gauge LNAH models with adjoint scalar fields is needed to achieve a satisfactory characterization of their main features.
@@ 9.5 Lattice SO($N_{c}$) gauge models with multiflavor scalar fields
One may again consider the unit-length limit of the scalar field, formally obtained by taking the limit $r,u\to\infty$ keeping $r/u=-1$ fixed, leading to the constraint ${\rm Tr}\,\varphi_{\bm{x}}^{t}\varphi_{\bm{x}}=1$. In the one-flavor case, $N_{f}=1$, the phase diagram of the lattice SO($N_{c}$) gauge model (117) is expected to show only one phase. This can be easily verified for $\gamma=0$, where the model becomes trivial in the unitary gauge, and cannot have any phase transition. For $N_{c}=2$ the gauge group is equivalent to the Abelian group U(1), and the model is equivalent to LAHMs with compact gauge variables discussed in Sec. 6. New critical behaviors are expected only in the non-Abelian case, $N_{c}\geq 3$. The natural order parameter for the breaking of the SO($N_{f}$) global symmetry is the gauge-invariant real traceless and symmetric bilinear operator
@@
The field-theoretical $\varepsilon$-expansion analysis of the RG flow of the NAHFT (97) predicts the existence of charged universality classes for a sufficiently large number $N_{f}$ of components, controlled by the stable charged FP appearing for positive values of $v$. No stable FPs emerge for $v<0$, at least close to four dimensions. Beside charged transitions statistical systems may also develop other critical behaviors described by appropriate LGW field theories, in which the gauge correlations are not critical, as is the case in the strong-coupling ($\gamma\to 0$) gauge limit.
[…]
@@
Note that charged continuous transitions are possible for large values of $N_{f}$ according to the RG flow of the SO($N_{c}$) gauge NAHFT, see Sec. 8.2, because of the presence of a stable large-$N_{f}$ charged FP [304, 408, 186]. As it occurs in SU($N_{c}$) gauge LNAH models with scalar fields in the fundamental representation, continuous charged transitions are expected to appear for positive values of $v$, sufficiently large values of $\gamma$, and a sufficiently large number $N_{f}$ of components.
10 Two-dimensional lattice Abelian and non-Abelian gauge Higgs models
@@
In the previous sections we discussed 3D Abelian and non-Abelian Higgs models, to investigate the role of the gauge and global symmetries in determining their main thermodynamic properties, such as the Higgs phases, the nature of the phase transitions, and the continuum limits associated with their critical behaviors. Analogous issues can be addressed in lower-dimensional models.
@@
In two dimensions the phase diagram of lattice models with short-range interactions (such as the nearest-neighbor models considered in the previous sections) and continuous global symmetries are generally characterized by a single disordered phase. Indeed, according to the Mermin-Wagner theorem [454, 455], they cannot have magnetized phases characterized by the condensation of an order parameter, and therefore they do not undergo phase transitions driven by the spontaneous breaking of a continuous global symmetry.
@@
However, in the zero-temperature limit they can develop an asymptotic critical behavior, whose universal features are determined by the symmetries of the system and are generally associated with 2D nonlinear $\sigma$ models defined on symmetric spaces, see, e.g., Refs. [456, 2, 96]. The O($N$) $\sigma$ model with $N\geq 3$ and the CPN-1 model with $N\geq 2$ are paradigmatic models that show this type of behavior. Indeed, the lattice O($N$) model behaves as the nonlinear $\sigma$ model defined on the O($N$)/O($N-1$) symmetric space, while the CPN-1 lattice model is related to the nonlinear $\sigma$ model defined on the U($N$)/[U(1)$\times$×U($N-1$)] symmetric space, see, e.g., Ref. [2]. In this type of models, the correlation functions in the thermodynamic limit are characterized by a length scale $\xi$ that diverges as $T^{p}e^{c/T}$ for $T\to 0$, where $p$ and $c$ are universal coefficients that only depend on the universality class and that can be determined in perturbation theory (the coefficient $c$ also depends on the normalization of $T$, which can be uniquely fixed perturbatively), see A.5 for details.
@@
In this section we discuss how the interplay of global and gauge symmetries determines the effective low-energy field theory realized in the zero-temperature continuum limit (critical behavior) of 2D lattice gauge models using the standard Wilson formulation of gauge variables [101, 3].
@@ 10.1 General conjecture for the zero-temperature continuum limit
In the case of models characterized by both global and gauge symmetries, the asymptotic zero-temperature critical behavior (equivalently, continuum limit) is expected to arise from the interplay between the two different symmetries. For the purpose of understanding which features are relevant and which continuum limits are effectively realized, several 2D lattice Abelian and non-Abelian gauge models have been investigated [460, 461, 250, 462, 453, 463, 464]. The results of these analyses have led to the following general conjecture:
@@
2D field-theoretical models defined on symmetric spaces have been largely investigated, see, e.g., Ref. [2], and several high-order perturbative computations, in particular of their $\beta$ functions, have been reported in the literature [465, 466, 467, 468, 469, 470, 311, 456, 304, 471, 472, 473, 474, 475, 476, 477, 478, 479, 480, 2, 96]. The O($N$) $\sigma$ and the CPN-1 models are the simplest examples. Other examples of lattice gauge models with critical behavior corresponding to a symmetric-space field theory will be discussed in the following of this section.
10.2 Continuum limits of multiflavor lattice gauge theories
We now summarize the main results obtained for several 2D multiflavor lattice Abelian and non-Abelian gauge models [460, 461, 462, 453, 463, 464]. They support the conjecture on the nature of their continuum limit presented in Sec. 10.1. The analysis of the zero-temperature behavior requires the determination of the relevant low-energy degrees of freedom, which correspond to the minimum configurations of the Hamiltonian, and of the residual global symmetry, which essentially determines the appropriate corresponding field-theoretical $\sigma$ model. These analyses are then supplemented by nonperturbative FSS analyses of MC data.
10.2.1 Lattice U(1) gauge models
@@
For some particular values of $\gamma$ the model (120) becomes equivalent to simpler well-known models. In the limit $\gamma\to\infty$, the gauge link variables are all equal to one modulo gauge transformations, so model (120) becomes equivalent to the standard 2D O($2N$)-symmetric vector model, whose properties are reviewed in A.5. For $\gamma=0$ the AH model (120) is instead a particular lattice formulation of the CPN-1 model [2, 233, 234, 235, 475, 476].5151512D CPN-1 models have been much studied in the literature, both analytically and numerically, because they provide a theoretical laboratory to understand some of the mechanisms of quantum field theories of fundamental interactions. In particular, they share some notable features with QCD, the theory of the hadronic strong interactions, such as the asymptotic freedom and the so-called $\theta$ dependence related to topology, see, e.g., Refs. [469, 481, 470, 474, 482, 483]. For both $\gamma=0$ and $\gamma=\infty$, the lattice model shows a universal critical behavior for $\beta\to\infty$. In both cases, the correlation length $\xi$ increases exponentially, $\xi\sim\beta^{-p}e^{c\beta}$, with $p$ and $c$ taking different values, as in the CPN-1 and O($2N$) $\sigma$ models, see A.5.
As in the 3D LAH models discussed in Sec. 6, the zero-temperature critical behavior can be investigated by analyzing the correlations of the gauge-invariant bilinear $Q_{\bm{x}}^{ab}=\bar{z}_{\bm{x}}^{a}z_{\bm{x}}^{b}-\delta^{ab}/N$. The FSS analyses of the MC results reported in Ref. [460] show that the asymptotic behavior is independent of $\gamma$, and corresponds to that found for $\gamma=0$. Therefore, the zero-temperature critical behavior of the 2D lattice U(1) model with an $N$-component complex scalar field belongs to the universality class of the 2D CPN-1 model, thus corresponding to the symmetric space U($N$)/[U(1)$\times$×U($N-1$)]. The independence of the critical behavior on $\gamma$ is related to the subleading behavior of the gauge modes.
10.2.2 Lattice SU($N_{c}$) gauge models
We now consider 2D non-Abelian SU($N_{c}$) models defined on square lattices, with multiflavor matter fields in the fundamental and adjoint representations, analogous to those defined in Secs. 9.1 and 9.4.1, respectively.
@@ 10.2.2 Lattice SU($N_{c}$) gauge models
Figure 24: (Adapted from Ref. [461]) We show numerical results for the 2D SU($N_{c}$) gauge LNAH models with $N_{c}=3$ (left) and $N_{c}=4$ (right); $N_{f}=2$ and $\gamma=0$ in both cases. We plot the Binder parameter $U_{P}$ associated with the operator $P_{\bm{x}}^{fg}$, defined in Eq. (108), versus $\xi_{P}/L$, where $\xi_{P}$ is the second-moment correlation length associated with the correlation function $G_{P}$, defined as in Eq. 150. In both panels, with increasing $L$, data approach the solid line representing the universal scaling curve of the 2D CP1 model (see Ref. [460] for a parameterization), associated with the symmetric space that has the same global symmetry, thus supporting the conjecture reported in Sec. 10.1. As discussed in Sec. A.3.4), this comparison does not involve any nonuniversal normalization.
10.2.3 Lattice SO($N_{c}$) gauge models